# Thermalization and irreversibility of an isolated quantum system

### v1

@@ Abstract
Summary and outlook.— From the perspective of quantum information, we demonstrate the mechanism of irreversibility and thermalization in the $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence through local erasure of spreaded nonequilibrium initial state information. This mechanism is closely related to research on the decoherence[30], eigenstate thermalization hypothesis, quantum information scrambling[31, 32, 33, 34], and information erasure[35, 36, 37, 38], and further theoretical investigations are necessary. This mechanism has significant relevance for current research in quantum computing and quantum simulation technologies. In current quantum computing chips, Hamiltonians similar to Eq. (Thermalization and irreversibility of an isolated quantum system) are present, and unavoidable local perturbations can affect the storage and process of quantum information. Furthermore, our work demonstrates that the mechanisms of quantum information erasure caused thermalization can be simulated and studied using superconducting quantum chips. Quantum simulation experiments may offer even richer insights into these phenomena.

### v2

@@ Abstract
Summary and outlook.— From the perspective of quantum information, we demonstrate the mechanism of irreversibility and thermalization in the $[\hat{O}^{\dagger}\hat{O}(t)]^{N}$ sequence through local erasure of spread non-equilibrium initial state information. This mechanism is closely related to research on the decoherence[30], eigenstate thermalization hypothesis, quantum information scrambling[31, 32, 33, 34], and information erasure[35, 36, 37, 38], and further theoretical investigations are necessary. Future quantum simulations will enable unprecedented exploration of these phenomena.
Appendix A Supplemental Material
A.1 Quantum simulation on a superconducting quantum processor
Quantum simulations enable the study of the quantum dynamical evolution of the Hubbard model. Using the model in Fig. 4(a) as an example, we illustrate how such research can be conducted on a superconducting quantum chip. A 14-qubit chip can be constructed, as shown in Fig. 4(c), consisting of two 6-qubit rings forming inner and outer circles, with two additional qubits connected to the outer ring. Between neighboring qubits within each ring, there is a nearest-neighbor XY interaction, while between corresponding qubits in the inner and outer rings, there is a ZZ interaction. The other two qubits have only ZZ coupling with the qubits in the outer ring. The $XY$ interaction term corresponds to the hopping between neighboring lattice sites in our model, while the $ZZ$ interaction represents the repulsive or attractive interaction between different particles at the same site. Interaction parameters between qubits on the superconducting qubit chip can be tuned via couplers. Additionally, the coupling strength and sign between superconducting qubits can be controlled through these couplers, enabling quantum simulation of the $\hat{O}^{\dagger}e^{i\hat{H}t}\hat{O}e^{-i\hat{H}t}$ sequence evolution. There have been some experimental demonstrations of dynamical evolution in one-dimensional bosonic Hubbard models on superconducting qubit chips [26], as well as studies on thermalization and scrambling in bosonic Hubbard models on ladder structures [24, 23], and other experiments on two-dimensional bosonic Hubbard models [27, 28, 29].